a piece of ice slides down a 45degree incline in twice the time it takes to slide down a frictionless 45degree incline. find the coefficient of kinetic friction between ice and incline.


the answer is 0.75 but i want the solution the answer is given in the back side of my book . i.e. yugbodh physics class 11th.
1
Expert's answer
2012-09-28T08:42:59-0400
If the time is doubled by the presence of friction, the acceleration
is
reduced by a factor 4. (That is because S = a * t^2 / 2 remains the
same).
The friction is therefore 3/4 of the component of the weight
down the
incline.
m * g * cos 45 * k = (3/4) * m * g * sin 45, where k is the
kinetic
friction coefficient.
k = 3/4 = 0.75